# Breaking Focus: Contextual Distraction Curse in Large Language Models

### v1

@@ 3.5 Efficiency Strategies
where $[1:n_{1}]$ indicates the top $n_{1}$ nodes according to their values $\mathcal{V}(P^{\prime}_{j})$. The branch corresponding to $P^{\prime}$ is then terminated.
@@
Problem Filtering via Classifier. To reduce search costs, a classifier $C(Q)$ is used to filter out questions with low potential to become effective problem candidates (e.g., the extremely easy question “What is the highest mountain in the world?”). The classifier is trained on previously searched questions, $\mathcal{D}_{\text{train}}=\{(Q_{i},y_{i})\}_{i=1}^{N}$, where $y_{i}\in\{0,1\}$ indicates whether $Q_{i}$ successfully exposes a vulnerability in the victim model $M$.
@@
We show the overall algorithm in Algorithm 1.
Algorithm 1 Overall Algorithm
Output: Candidate problem list $L$
Initialize priority queue $\mathcal{Q}\leftarrow\emptyset$ and candidate list $L\leftarrow\emptyset$
foreach $P=\langle Q,A_{\text{gt}},\mathcal{A}_{\text{inc}}\rangle\in D$ do
       if $p(y=1\mid Q)=C(Q)<\tau_{C}$ then
             continue // Filter low-potential questions using classifier
       end if
      if $r_{M}(P)$ = 0 then
             Add $P$ to $L$continue
       end if
        Add root node $P$ to $\mathcal{Q}$
end foreach
while $\mathcal{Q}\neq\emptyset$ do
       Pop $P^{\prime}=\arg\max_{P\in\mathcal{Q}}\mathcal{V}(P)$, $\mathcal{Q}\leftarrow\mathcal{Q}\setminus\{P^{\prime}\}$  Generate $k=|\mathcal{A}_{\text{inc}}|$ child nodes for $P^{\prime}$ using $P_{\text{proxy}}$for each child node $P^{\prime}_{j}$ do
             Compute semantic shift $S(P,P^{\prime}_{j})$ and length ratio $\text{len}(P^{\prime}_{j})/\text{len}(P)$ // Semantic shift check and computing length ratio if $S(P,P^{\prime}_{j})=1$ and $\text{len}(P^{\prime}_{j})/\text{len}(P)\leq\lambda$ then
                   Compute value $\mathcal{V}(P^{\prime}_{j})$
             end if
            else
                   Discard $P^{\prime}_{j}$
             end if
       end for
      if $|\mathcal{C}_{0}(P^{\prime})|>n_{1}$ then
             Add top $n_{1}$ nodes from $\mathcal{C}_{0}(P^{\prime})$ to $L$ terminate branch  continue
       end if
      if $r_{M}(P^{\prime}_{j})=1,\,\forall P^{\prime}_{j}$ then
             continue // Skip nodes where all children are unpromising
       end if
      Add $P^{\prime}_{j}$ nodes to $\mathcal{Q}$if $m_{i+1}>m_{i},\,\forall i\in\{1,2,\dots,l-1\}$ then
             continue // Prune monotonically increasing success rate branches
       end if
end while
return $L$
@@ 4.1 Experiment Setup
Selected Datasets. We selected four widely used datasets for CDV investigation: MMLU (Hendrycks et al., 2021b, a), CommonsenseQA (Talmor et al., 2019), OpenbookQA (Mihaylov et al., 2018), and TruthfulQA (Lin et al., 2021). These datasets are widely used to evaluate LLMs on various dimensions, including commonsense reasoning and elementary science knowledge.
Models. As shown in Table 6, we used four proprietary models in our experiments: GPT-4o (Hurst et al., 2024), GPT-4o-mini (OpenAI, 2024), Claude-3.5-Sonnet (Anthropic, 2024), and o1-mini (Jaech et al., 2024). Additionally, we included eight open-weight models: Gemma-2-2B, Gemma-2-27B (Team, 2024a), Qwen2.5-1.5B, Qwen2.5-7B, Qwen2.5-72B (Yang et al., 2024; Team, 2024b), Llama-3.1-8B (Meta, 2024b), Llama-3.1-70B (Meta, 2024a) and Phi-3.5-mini (Abdin et al., 2024).
Hyperparameter Setting. We set the temperature to 0.7 during the perturbation generation phase to encourage more creative and reproducible responses. For the evaluation phase, we lowered the temperature to 0.001 to ensure the stability of the responses, with a maximum output length of 1,024 tokens. Additionally, we set $\alpha$ to 2 and $\gamma$ to 1 in all experiments. For other detailed hyperparameter settings, please refer to Appendix A.2.
Prompt Template. We employ prompt-based approaches to perform the following tasks: generating perturbations, evaluating semantic similarity, model assessment (zero-shot + CoT), enhancing problems as a baseline, identifying hard-to-perturb problems, and mitigating CDV. The specific prompt templates are provided in Appendix D.
4.2 Main Results
We conducted extensive evaluation experiments and mitigation experiments on CDV. The detailed configurations of each experiment and their correspondence with all the figures can be found in the Appendix A.
Our method enables LLMs to autonomously generate CDV examples and effectively engage in self-challenging. In Figure 2, the same model was configured to serve as both the proxy and victim model during the process of CDV investigation, we can observe that all tested models exhibit a significant decline in accuracy on perturbed questions when compared to their performance on the original questions. For instance, GPT-4o-mini’s performance significantly drops by more than 40%.
Furthermore, it also reveals an intriguing pattern the vulnerabilities identified by a model itself tend to be more challenging than those discovered by other models. For instance, the GPT-4o-mini exhibits a significantly lower accuracy of 0.185 on perturbed questions generated through its self-identified vulnerabilities, compared to its performance of 0.235 on perturbations created by the more sophisticated GPT-4o. This indicates that the model itself is more aware of its weaknesses, aligning with the justifiability of previous works about self-alignment (Sun et al., 2024) and self-correction (Pan et al., 2023).
Table 1: Model performance on original and perturbed datasets, and the corresponding performance decline.
@@ 4.2 Main Results
All models are susceptible to challenges posed by CDV examples, regardless of their capabilities. Our findings reveal that CDV examples generated by stronger models effectively challenge weaker models, while those generated by weaker models also significantly impact stronger ones. As shown in Figure 2, despite their superior language, reasoning, and knowledge capabilities, the stronger models have not achieved complete dominance over smaller models in handling CDV examples. Even the most advanced LLMs [^1] We refer to the leaderboard at https://lmarena.ai/ for model performance comparisons., such as GPT-4o and Claude-3.5-Sonnet, achieve less than 50% accuracy on questions perturbed by Gemma-2-27B. This underscores a fundamental concern, showing that no model is fully immune to such vulnerability.
Figure 3: Accuracy on original and perturbed samples by different methods on the sampled dataset. Elaborated: Semantic-preserving length augmentation, Prompt-only: Auto-generated perturbations without search, DyVal2: Dynamic evaluation framework (Zhu et al., 2024). Lower scores indicate stronger CDV challenge effectiveness.
Model Original Elaborated Prompt-only DyVal2 Ours
@@
The extent of performance degradation varies significantly across tasks. For example, as shown in Table 1, the average performance drop on the TruthfulQA dataset is consistently smaller than that observed on OpenbookQA for all evaluated models. This trend holds true across models, indicating that model sensitivity against CDV varies in different tasks. Tasks that emphasize domain-specific knowledge like OpenbookQA, appear to expose model vulnerabilities more effectively than trustworthiness-related tasks like TruthfulQA.
Our method outperforms existing perturbation techniques by a significant margin. We consider two traditional perturbation approaches: length-based augmentation (Elaborated), which expands the original question’s length while preserving its core meaning, and simple auto-generated perturbations (Prompt-only), which uses a single prompt to add interference to the question. However, these methods lead to only minimal performance drops. As shown in Figure 3, even advanced methods like DyVal2 achieve only a 15.2% average accuracy reduction. In contrast, our proposed perturbation method introduces a dramatic 52.0% average accuracy decline, outperforming existing methods by 2 to 4 times. Notably, even highly capable models such as Claude-3.5-Sonnet exhibit a substantial 49.5% absolute accuracy drop under our perturbations, compared to just 10.0% under DyVal2. These results highlight the unique strength of our tree-based search framework in systematically targeting the model’s ability to discern and resist distraction across diverse architectures, rather than merely increasing input complexity.
@@ 4.3 Classifier: Filtering Hard-to-Perturb Problems
A critical challenge in CDV generation lies in distinguishing between hard-to-perturb problems (the questions are quite easy and naive for LLMs so that it’s highly impossible to be the CDV examples, like elementary arithmetic or factual questions) and easy-to-perturb problems (questions susceptible to semantic-preserving perturbations). Our analysis reveals that approximately 37% of computational resources are typically wasted on enhancing hard-to-perturb problems. To address this inefficiency, we develop classifiers that predict the perturbation possibility of each problem.
Table 2: The impact of the classifier on the perturbation success rate of the LLMs. The full model names are: GPT-4o, Gemma-2-27B, Llama-3.1-70B, and Qwen2-5-72B. The rows display the perturbation success rate with and without the classifier.
@@
Firstly, to validate the classifier’s generalization, it is crucial to examine whether hard-to-perturb questions exhibit nearly consistency across most LLMs. The confusion matrix in Figure 4 demonstrates significant overlap in perturbation difficulty across models: about 82% of problems are either perturbable or non-perturbable for both models in any pairwise comparison. This consistency across models enables the development of universal classifiers for the prediction of perturbation difficulty.
@@
We implement two classifier paradigms: (1) Prompt-based classifiers using zero-shot LLM judgments, and (2) Fine-tuned classifiers trained on 1,080 annotated samples with 120 held-out test cases. Notably, as shown in Table 2, classifiers trained on GPT-4o-mini-derived data generalize robustly to diverse LLMs, maintaining high precision even on stronger models. The evaluation focuses on precision-recall tradeoffs using the Fβ metric ($\beta=0.5$), prioritizing precision to minimize false positives in perturbable problem selection. The Fβ score is computed as:
$\text{F}_{\beta}=(1+\beta^{2})\times\frac{\text{Precision}\times\text{Recall}}{(\beta^{2}\times\text{Precision})+\text{Recall}}$ (15)
The results shown in Figure 5 and Table 7 reveal that fine-tuned classifiers substantially improve perturbation efficiency through precise problem selection. The fine-tuned classifiers achieve 83% precision in identifying perturbable problems, outperforming the best prompt-based approach (GPT-4o at 68%). We will discuss the details of computational cost saved by the trained classifier in subsection 4.4.
@@ 4.4 Ablation Study
Impact of the value function. We evaluate the effectiveness of the designed value function $\mathcal{V}(P^{\prime})$ and its components introduced in subsection 3.4 about guiding the search process. This function incorporates two key factors: the success rate $r_{M}(P^{\prime})$ of the victim model on the perturbed problem $P^{\prime}$, and a depth penalty $\text{depth}^{-\gamma}$, designed to balance exploration and computational efficiency.
To assess the individual contributions of the success rate $r_{M}(P^{\prime})$ and the depth penalty to the overall effectiveness of the perturbations, we randomly selected questions from four different datasets for testing and calculated the perturbation success rate (the percentage of the questions that achieves 0% accuracy (i.e., $r_{M}(P^{\prime})=0$)). The results are as follows: using the complete value function, the perturbation success rate was 59%; without the depth penalty term $\text{depth}^{-\gamma}$, the perturbation success rate dropped to 57%; and without $r_{M}(P^{\prime})$, the perturbation success rate further decreased to 53%. These results demonstrate that both the success rate term and the depth penalty term significantly enhance the effectiveness of the search process.
Cost saved by classifier. To evaluate the impact of incorporating the fine-tuned classifier into the pipeline, we experimented with comparing the performance with and without the classifier. We randomly selected questions from four datasets for this study. For the condition utilizing the classifier, we applied our fine-tuned classifier to filter and select 100 questions identified as easy-to-perturb. For the condition without the classifier, we directly sampled 100 questions at random without any prior filtering.
The experimental results are presented in Table 3. By effectively filtering out hard-to-perturb questions, the classifier allowed us to focus resources on those that were more susceptible to successful perturbation. This led to a significant increase in the perturbation success rate and minimized the costs associated with unsuccessful attempts (the perturbation success rate increase by 38.9% and the average cost decreased by 22%).
4.5 Mitigation of CDV
Notably, models that perform well on original questions but fail on perturbed ones show they have the necessary knowledge but are susceptible to contextual information. To address this, we explored both prompt-based (i.e., training-free) and training-based strategies to improve their performance on these challenging questions.
Prompt-based mitigation is unable to patch CDV. Since our perturbed questions introduce context-based perturbations while preserving the core question content, we explored whether explicit prompt instructions could help models focus on the essential information and filter out the added distractions. We tested this approach by modifying the original prompts to include specific guidance on identifying and focusing on the key question components, with detailed prompt templates provided in Appendix D. As shown in Table 4, the prompt-based approach did not yield significant improvements in mitigating CDV. Some models, such as o1-mini and Claude-3.5-Sonnet, even showed slightly decreased accuracy with the modified prompts. This suggests that the distractions introduced by our perturbed questions cannot be effectively addressed through simple prompt refinements alone.
Table 3: Comparison of cost with and without classifier. Inp. Tok.: Number of input tokens, Out. Tok.: Number of output tokens, Pert. Ques.: Number of successfully perturbed questions, Pert. Rate (%): Perturbation success rate, Cost/Ques. (\$): USD Cost per perturbed question.
@@ 4.5 Mitigation of CDV
w/ classifier 3.81M 1.57M 236 82% ($\uparrow$ 38.9%) 0.0064 ($\downarrow$ 21.9%)
Training-based mitigation brings significant benefits. The limited success of prompt-based mitigation inspired us to explore whether training-based approaches (i.e., fine-tuning) could work well. We leveraged about 1,200 preference data pairs curated from our dataset, where each pair consisted of a question, a chosen answer (correct answer), and a rejected answer (wrong answer). These pairs were derived from model responses in our main experiments and were split into training and test sets. We selected three open-weight models for training, and the training process was conducted using the DPO (Rafailov et al., 2024).
Table 4: Model accuracy before and after prompt-based mitigation (Original vs Enh.).
@@
GPT-4o-mini 0.185 0.211 $+$0.026
Llama-3.1-8B 0.247 0.251 $+$0.003
Gemma-2-27B 0.237 0.255 $+$0.018
o1-mini 0.374 0.366 --0.008
Qwen2.5-72B 0.334 0.343 $+$0.009
GPT-4o 0.390 0.391 $+$0.002
Claude-3.5-sonnet 0.495 0.481 --0.013
As shown in Table 5, the results indicate that the performance of the original models was comparable to GPT-4o-mini, while significantly behind more powerful models like GPT-4o and Qwen2.5-72B. However, after fine-tuning, all three models demonstrated substantial improvements in accuracy on the enhanced questions. Most remarkably, Phi-3.5-mini achieved exceptional progress, surpassing all comparison models, including GPT-4o. Through detailed case studies provided in Appendix C, we found that fine-tuned models exhibited markedly improved ability to mitigate CDV rather than simply gaining new knowledge. Notably, a large fraction of the originally incorrect answers remained incorrect after training (e.g., 82.1% for Phi-3.5-mini), suggesting that the improvements were driven by a stronger ability to identify and ignore irrelevant contextual information, maintaining focus on the key points required to answer questions correctly.
Table 5: Model accuracy before and after DPO training. Retain: fraction of original incorrect answers that remain incorrect after training.
@@
GPT-4o 0.568 - - -
Qwen2.5-72B 0.519 - - -
GPT-4o-mini 0.232 - - -
These findings shed light on the nature of CDV and strategies for its mitigation. The limited effectiveness of prompt-based mitigation and significant improvement brought by training indicates that the interference caused by our enhanced questions impacts the models at a deeper decision-making level, rather than merely introducing superficial distractions (i.e., LLMs are unable to be aware of such distractions simply guided by prompt engineering). This highlights the importance of incorporating CDV mitigation into the initial training pipeline, and we urge the community to prioritize these aspects during model development, rather than relying on post-hoc solutions, which may be less effective and computationally expensive.

### v2

@@ 2.5 Efficiency Strategies
where $1:n_{1}]$ indicates the top $n_{1}$ nodes according to their values $\mathcal{V}(P^{\prime}_{j})$. The branch corresponding to $P^{\prime}$ is then terminated.
@@
Problem Filtering via Classifier. To reduce search costs, a classifier $C(Q)$ is used to filter out questions with low potential to become effective distraction candidates (e.g., the extremely easy question “What is the highest mountain in the world?”). The classifier is trained on previously searched questions, $\mathcal{D}_{\text{train}}=\{(Q_{i},y_{i})\}_{i=1}^{N}$, where $y_{i}\in\{0,1\}$ indicates whether $Q_{i}$ successfully exposes a vulnerability in the victim model $M$.
@@
Due to the space limitation, we show the overall algorithm in Appendix D.
@@ 3.1 Experiment Setup
Selected Datasets. We selected four widely used benchmarks to evaluate contextual robustness under adaptive distraction: MMLU (hendryckstest2021, ; hendrycks2021ethics, ), CommonsenseQA (talmor-etal-2019-commonsenseqa, ), OpenbookQA (OpenBookQA2018, ), and TruthfulQA (lin2021truthfulqa, ). These datasets cover diverse domains such as factual knowledge, commonsense reasoning, and elementary science, making them suitable for probing how LLMs handle irrelevant but semantically coherent contextual additions.
Models. As shown in Table 7, we used four proprietary models in our experiments: GPT-4o (hurst2024gpt, ), GPT-4o-mini (openai2024gpt4omini, ), Claude-3.5-Sonnet (anthropic2024claude35, ), and o1-mini (jaech2024openai, ). Additionally, we included eight open-weight models: Gemma-2-2B, Gemma-2-27B (gemma_2024, ), Qwen2.5-1.5B, Qwen2.5-7B, Qwen2.5-72B (qwen2, ; qwen2.5, ), Llama-3.1-8B (meta2024llama31_8b, ), Llama-3.1-70B (meta2024llama31_70b, ), and Phi-3.5-mini (abdin2024phi, ).
Hyperparameter Setting. We set the temperature to 0.7 during the distraction generation phase to encourage more diverse and challenging outputs. For evaluation, we lowered the temperature to 0.001 to ensure response consistency, with a maximum output length of 1,024 tokens. Additionally, we set $\alpha=2$ and $\gamma=1$ for the value function used in the tree search. For other detailed hyperparameter settings, please refer to Appendix B.1.
Prompt Template. Prompt-based templates are used for several sub-tasks throughout our framework, including generating contextual distractions, assessing semantic consistency, evaluating model answers (zero-shot + CoT), baseline elaboration, filtering distraction-susceptible samples, and conducting robustness enhancement. The specific templates are provided in Appendix G.
Human Verification. To confirm semantic preservation, we conducted a human evaluation on randomly sampled perturbed questions. Annotators judged both semantic equivalence and answer consistency. Detailed results are reported in Appendix C.
Table 1: Accuracy of seven LLMs on four benchmarks before (Original) and after (Perturbed) applying our adaptive distractions. Cell background colors emphasise the severity of that drop.
@@
3.2 Main Results
We conducted extensive evaluation experiments and mitigation studies to assess the impact of adaptive distractions on LLMs. The detailed configurations and dataset-model setups corresponding to each figure and table can be found in Appendix B.
Our method enables LLMs to autonomously generate adaptive distractions and effectively self-challenge. In Figure 2, we configure the same model to act as both the proxy and the victim in the distraction generation process. We observe that all tested models suffer a substantial drop in accuracy when evaluated on adaptively distracted questions, compared to their original performance. For instance, GPT-4o-mini experiences a performance decline of over 40%.
Furthermore, we uncover an intriguing pattern: distractions generated by a model itself tend to be more adversarial to that model than those generated by others. For example, GPT-4o-mini achieves an accuracy of only 0.185 on distractions it generated for itself, compared to 0.235 on those generated by the more advanced GPT-4o. This suggests that models may be better at identifying their own weaknesses, resonating with previous findings on self-alignment sun2024principle and self-correction pan2023automatically .
All models are susceptible to adaptive distractions, regardless of their scale or capability. Our results show that distractions created by stronger models can reliably challenge weaker models, while even distractions from weaker models can degrade the performance of stronger ones. As shown in Figure 2, powerful models such as GPT-4o and Claude-3.5-Sonnet still achieve below 50% accuracy when evaluated on distractions generated by Gemma-2-27B. This highlights a fundamental vulnerability: no model is currently robust to adaptive distractions. [^1] We refer to the leaderboard at https://lmarena.ai/ for model performance comparisons.
The extent of performance degradation varies significantly across tasks. As shown in Table 1, the average drop in accuracy on TruthfulQA is consistently smaller than that on OpenbookQA across all models. This suggests that distraction sensitivity varies by domain. Tasks like OpenbookQA, which require precise factual retrieval, appear more vulnerable to contextual interference than trustworthiness-based tasks such as TruthfulQA.
Table 2: Accuracy on original and distracted samples generated by different methods. ICA: Adding semantically coherent but task-irrelevant information to increase question length. SPD: Generating distractions via a single prompt without further optimization. DyVal2: Dynamic evaluation baseline (zhu2024dynamic, ). Lower scores indicate more effective distractions.
Model Original ICA SPD DyVal2 Ours
@@ 3.2 Main Results
Our method outperforms existing distraction techniques by a large margin. We compare our method with several baselines: (1) Irrelevant Context Augmentation (ICA), which adds semantically coherent but task-irrelevant information to extend question length ye2024justice ; (2) Single-Prompt Distraction (SPD), which generates distractions via a single prompt without optimization (shi2023large, ); and (3) DyVal2, a recent dynamic evaluation framework (zhu2024dynamic, ). As shown in Table 2, our method results in an average accuracy drop of 52.0%, compared to 15.2% for DyVal2. Even strong models such as Claude-3.5-Sonnet experience a 38.4% absolute accuracy drop under our framework, compared to only 10.0% under DyVal2. These results highlight the strength of our tree-based search framework in systematically identifying contextual vulnerabilities, not merely increasing input complexity.
Adaptive distractions demonstrate robust cross-task generalization. We further examine the generalization of adaptive distraction by applying it to other tasks with well-defined ground truth, including mathematical reasoning. As shown in Appendix B.6, our method remains effective on the MATH-500 benchmark(lightman2023lets, ), suggesting that contextual distraction is not confined to factual QA but extends to any task where correctness can be explicitly evaluated.
@@ 3.3 Classifier: Filtering Hard-to-Perturb Problems
A critical challenge in adaptive distraction generation lies in distinguishing between hard-to-perturb problems (e.g., simple factual or arithmetic questions that are consistently answered correctly by LLMs regardless of added context) and easy-to-perturb problems (questions susceptible to semantic-preserving contextual distractions). Our analysis reveals that approximately 37% of computational resources are typically wasted on attempting to distract hard-to-perturb examples. To address this inefficiency, we design classifiers that predict the distraction susceptibility of a given question.
To evaluate the generalizability of the classifier, we examine whether perturbation difficulty is consistent across different LLMs. As shown in the confusion matrix in Figure 4, there is strong alignment between models: around 82% of questions are either distractable or non-distractable for both models in any pairwise comparison. This consistency suggests that model-agnostic classifiers can be trained to identify distractable inputs.
We implement two types of classifiers: (1) Prompt-based classifiers, which leverage LLM Judge (ye2024justice, ; wang2025trusteval, ), and (2) Fine-tuned classifiers, trained on 1,080 annotated examples with 120 held-out test cases. As shown in Table 3, classifiers trained on data from GPT-4o-mini generalize effectively to stronger models, maintaining high precision across model families. We use the Fβ score with $\beta=0.5$ to prioritize precision. The formal definition is provided in Appendix B.1.
As illustrated in Figure 3 and Table 8, fine-tuned classifiers significantly improve overall efficiency by accurately filtering out hard-to-perturb samples. In particular, our best fine-tuned classifiers achieve up to 83% precision on identifying distractable problems, outperforming the best prompt-based baseline (GPT-4o) at 68%. The reduction in wasted computational effort and improvement in overall perturbation success rates will be analyzed in more detail in Experiment 3.4.
To further assess whether the classifier generalizes across tasks, we conduct a cross-dataset evaluation. Specifically, we train the classifier using three datasets (MMLU, CommonsenseQA, and OpenbookQA), and validate its performance on the held-out TruthfulQA. Table 10 presents the $F_{0.5}$ scores under this setting. Despite not being trained on TruthfulQA, the classifier maintains comparable performance, suggesting that it captures general signals of distraction susceptibility rather than dataset-specific patterns.
Table 3: The impact of the classifier on the perturbation success rate of the LLMs. The full model names are: GPT-4o, Gemma-2-27B, Llama-3.1-70B, and Qwen2.5-72B. The rows display the perturbation success rate with and without the classifier.
@@ 3.4 Ablation Study
Impact of the value function. We evaluate the effectiveness of the designed value function $\mathcal{V}(P^{\prime})$ and its components introduced in Method 2.4 for guiding the search process. This function incorporates two key factors: the success rate $r_{M}(P^{\prime})$ of the victim model on the distracted version of the problem $P^{\prime}$, and a depth penalty $\text{depth}^{-\gamma}$, which helps balance the trade-off between exploration and computational efficiency.
To assess the individual contributions of these components, we randomly selected questions from four datasets and measured the distraction success rate (defined as the percentage of questions for which the model fails completely under distraction, i.e., $r_{M}(P^{\prime})=0$). The results show that using the complete value function yields a 59% success rate. Removing the depth penalty reduces this rate to 57%, and removing the success rate term further decreases it to 53%. These findings confirm that both the model failure signal and the search depth control play important roles in identifying effective adaptive distractions.
Cost saved by classifier. To examine the impact of the classifier, we compared performance with and without classifier filtering. In this experiment, we randomly selected questions from four datasets. In the classifier condition, we used our classifier to select 100 questions predicted to be susceptible to distraction. In the baseline condition, we sampled 100 questions at random without any filtering.
Table 4: Computational cost comparison with and without classifier. Inp. Tok.: Number of input tokens, Out. Tok.: Number of output tokens, Pert. Ques.: Number of successfully distracted questions, Pert. Rate (%): Distraction success rate, Cost/Ques. (\$): USD cost per successfully distracted question.
@@
[…]
As shown in Table 4, incorporating the classifier significantly improves both the effectiveness and the efficiency of the generation process. By filtering out hard-to-distract questions, the classifier enables better allocation of resources to more promising inputs. This leads to a 38.9% increase in success rate and a 21.9% reduction in average cost per successful distraction.
3.5 Mitigation under Adaptive Distraction
Notably, models that perform well on original questions but fail on distracted ones demonstrate that they possess the necessary knowledge yet remain vulnerable to contextual interference. To address this issue, we explored both prompt-based (i.e., training-free) and training-based strategies to improve their performance on these challenging questions.
Table 5: Model accuracy before and after prompt-based mitigation (Orig. vs Enh.).
@@ 3.5 Mitigation under Adaptive Distraction
GPT-4o-mini 0.185 0.211 ++0.026
Llama-3.1-8B 0.247 0.251 ++0.003
Gemma-2-27B 0.237 0.255 ++0.018
o1-mini 0.374 0.366 -0.008
Qwen2.5-72B 0.334 0.343 ++0.009
GPT-4o 0.390 0.391 ++0.002
Claude-3.5-sonnet 0.495 0.481 -0.013
Prompt-based mitigation is unable to alleviate distraction vulnerability. Since our adaptive distractions preserve the core question semantics while introducing task-irrelevant context, we tested whether adding explicit instructions in the prompt could help models focus on the essential content and ignore misleading information. Specifically, we modified the original prompts to include guidance on identifying and prioritizing key components of the question. Detailed templates are provided in Appendix G. As shown in Table 5, this approach yielded only marginal improvements. Some models, such as o1-mini and Claude-3.5-Sonnet, even showed slightly lower accuracy after prompt modifications. This suggests that the contextual interference introduced by adaptive distractions cannot be effectively mitigated through prompt refinement alone.
We also tested additional prompting methods that have shown promise in prior work, including In-context learning wei2022chain and Self-consistency wang2022self . As shown in Appendix B.5, none of these approaches substantially recover the original performance when faced with adaptive distractions. This reinforces our finding that prompt-based methods alone offer limited robustness.
Table 6: Model accuracy before and after DPO training. Retain shows the fraction of original incorrect answers that remain incorrect after training.
@@
GPT-4o 0.568
Qwen2.5-72B 0.519
GPT-4o-mini 0.232
As shown in Table 6, all three models benefited significantly from training. The Phi-3.5-mini model, in particular, achieved a post-training accuracy that surpassed even GPT-4o on the same distracted inputs. Detailed case studies in Appendix F show that the improvements were not merely due to new knowledge acquisition. Rather, fine-tuned models showed better focus on relevant question content and stronger resistance to irrelevant distractions. A large fraction of the originally incorrect answers remained incorrect, such as 82.1% for Phi-3.5-mini, suggesting that performance gains came from improved robustness rather than memorization. Additional analysis in Appendix B.7 confirms that DPO offers greater gains than supervised fine-tuning, while preserving performance on clean questions, which further validates that our improvements stem from true robustness rather than memorization.